OpenStudy (anonymous):

how do you evaluate the expression -3g - 2 -(-4g + 10) when g = -6 ?

OpenStudy (kropot72):

Substitute the value for g in the expression as follows: (-3*-6) - 2 -((-4*-6) + 10)) = -18 -2 -(24 + 10) = -20 -34 + -54

OpenStudy (anonymous):

so the answer is -18?

OpenStudy (kropot72):

Sorry the final result has a typo. It should read -20 -34 = -54